Water is added to the solution such that the mole fraction of water in the solution becomes 0.9. The boiling point of the solution is 
a.354.7K        b.376.2K

Grade:

3 Answers

Arun
4 years ago
 
I have attached an image containing solution. please check and let me know if any query exist. I will be happy to help you.
Rajat
4 years ago
The actual boiling point of water when no solute is added in it is 373 K. When a solute is added to an aqueous solution according to the colligative property the solution suffers a increase in boiling point.
Thus here the boiling point of the solution must be greater than 373 K. So answer is B.
Increase in boiling point =∆T= K(m/W)
Arun
4 years ago
Dear student
X(H2O) = 0.9
X (C2H5OH) = 0.1
delta Tb = Kb * m
0.52 * 0.1 * 1000/ (0.9 * 18) = 3.21 K
hence boiling point Tb = 373 + 3.21 = 376.21K
 
Hence option B is correct